OpenStudy (anonymous):

(1-cos2x)(1-cos2x)(1-cos2x) =

OpenStudy (loser66):

a * a *a = a^3 , right?

OpenStudy (aum):

\[ 1 - \cos(2x) = 2\sin^2(x) \]